An oil wildcatter has rights to a piece of land. If she drills, there is a 20% chance that she will strike oil, in which case the payoff is $200,000. The cost of drilling is $20,000. a) Should she drill? b) What is the most you would pay for the services of a geologist who can assess the oil prospects with perfect accuracy?
Answer #1

a)

here as expected value of drilling =0.2*200000=40000

as expected value is higher then cost of drilling therefore one should drill

b)

maximum amount paid for services =40000-20000=20000